OpenStudy (anonymous):

Help with algebra 2 homework!!!

OpenStudy (anonymous):

Hello?

OpenStudy (across):

What do you want us to do? Do you want us to read your mind? You have to ask a question!

OpenStudy (anonymous):

Sorry, D=kA[(t-y)/L] solving for y

OpenStudy (across):

You have that\[D=kA\frac{t-y}{L}\]and are asked to solve for \(y\). Is that correct?

OpenStudy (anonymous):

Yes I am solving for y, but there are [ ] around the fraction

OpenStudy (across):

What have you tried already? Also, the square brackets are optional.

OpenStudy (anonymous):

I don't understand how to get rid of the square brackets because I've learned that those mean the answer within the square brackets mean that it will always turn out to be positive

OpenStudy (across):

Oh, in that case, you meant to express\[D=kA\left|\frac{t-y}{L}\right|\]Is that right?

OpenStudy (anonymous):

Yes

OpenStudy (across):

You have to know that\[kA\left|\frac{t-y}{L}\right|\neq kA\left[\frac{t-y}{L}\right]\]

OpenStudy (across):

The left hand side means "absolute value."

OpenStudy (anonymous):

It seems to have the brackets and not the absolute value sign aeound it

OpenStudy (across):

In that case, I assume that you must have gotten this far:\[\begin{align}D&=kA\left|\frac{t-y}{L}\right|\\\frac D{kA}&=\left|\frac{t-y}{L}\right|\end{align}\]

OpenStudy (across):

If it has square brackets, then you can solve it by simple multiplication and addition.

OpenStudy (anonymous):

I'm sorry, I thought that the brackets meant absolute value, and if they didn't, don't you have to distribute

OpenStudy (across):

You could, but it's not necessary.

OpenStudy (anonymous):

Alright, thank you so much